# A Scalable Pipeline for Enabling Non-Verbal Speech Generation and Understanding

### v1

@@ 5 Experiments
Baseline Models
CosyVoice2 0.090 0.074 3.930 5.023 3.240 3.288 0.514 0.671 4.09 4.06 2.29 2.54
Dia 0.182 0.097 15.687 171.882 2.749 2.528 0.431 0.319 3.33 3.51
F5-TTS and Variants
F5-TTS 0.044 0.036 5.325 6.188 3.046 3.200 0.673 0.721 3.61 3.22
   + CapSpeech 0.147 0.119 10.539 16.137 2.889 3.048 0.563 0.661 2.75 2.01 2.02 2.11
   + NVTTS 0.103 0.076 8.013 15.808 3.069 3.149 0.587 0.648 3.63 2.53 3.25 2.80
   + NVS (Ours) 0.104 0.102 10.132 7.341 3.033 3.115 0.544 0.671 3.43 3.63 3.22 3.44
0.137 0.146
Table 4: Evaluation results of non-verbal speech generation on English (EN) and Chinese (ZH) datasets. The best and second-best values in each column are highlighted in bold and underline, respectively. CLAP Score and IMOS are computed only on tags supported by each model or dataset, as specified in Table 5. The last row presents the CLAP Score of F5-TTS fine-tuned on our proposed dataset, evaluated solely on tags supported by CapSpeech. Here, NVTTS stands for NonVerbalTTS, and NVS stands for NonVerbalSpeech-38K.
@@ 5.1 Experimental Setup for Non-Verbal Speech Generation
In the non-verbal speech generation task, we use F5-TTS (Chen et al. 2024) as the base model. To support non-verbal control, we expand the vocabulary of F5-TTS by adding non-verbal tags such as [laughing], [coughing]. We fine-tune F5-TTS on three datasets: CapSpeech (Wang et al. 2025), NonVerbalTTS (Borisov, Spirin, and Diatlova 2025), and our proposed NonVerbalSpeech-38K dataset. More detailed information can be found in the Appendix.
@@
In addition to fine-tuning F5-TTS on different datasets, we compare it with two strong baselines: CosyVoice2 (Du et al. 2024b) and Dia (Nari-labs 2025), both of which support non-verbal control and generation (Table 5).
@@
Due to differences in the non-verbal tags supported by each model and dataset (Table 5), we select six common non-verbal phenomena for evaluation: [laughing], [sniff], [breath], [coughing], [throatclearing], and [sigh]. To evaluate the non-verbal speech generation capabilities of different models, we reuse the prompt speech provided by the seed-tts-eval framework (Anastassiou et al. 2024). For the text inputs, we generate tagged sentences using Genimi (Team et al. 2023), GPT-4o (Hurst et al. 2024). For each tag, we construct 175 sentences in both Chinese and English, resulting in a total of 1,050 sentences per language.
Label Datasets Models
Cap NVTTS NVS38K CV2 Dia
Coughing Y Y Y Y Y
Laughing Y Y Y Y Y
Sigh N Y Y N Y
Sniff N Y Y N Y
Throat Clearing N Y Y N Y
Breath Y Y Y Y N
Table 5: Non-verbal label support across datasets and models. Cap: CapSpeech, NVTTS: NonVerbalTTS, NVS38K: NonVerbalSpeech-38K, CV2: CosyVoice2.
To evaluate system performance, we use standard metrics including DNSMOS, SSIM, and WER/CER to assess speech quality, speaker similarity, and intelligibility. Additionally, we introduce the CLAP Score to measure the presence of intended non-verbal phenomena in synthesized speech. For subjective evaluation, 20 participants rated the samples on QMOS (for overall speech quality) and IMOS (for accuracy in rendering both verbal content and non-verbal tags). See Appendix for scoring details.
Model WER Details F1-Score $\uparrow$ TPD $\downarrow$ NTD $\downarrow$
WER (%) WER << 50% (%) # WER << 50%
Whisper-Large-V3 17.542 8.693 731 / 754
Qwen2-Audio 33.087 11.603 620 / 754
   + CapSpeech 23.806 16.508 709 / 754 0.598 4.232 0.172
   + NonVerbalSpeech-38K (Ours) 20.215 14.445 727 / 754 0.661 6.223 0.284
Table 6: Evaluation results for end-to-end non-verbal speech understanding. WER << 50% refers to WER calculated after removing samples with extremely high error rates (WER >> 50%). # WER << 50% denotes the count of samples with WER below 50% over the total number of test samples.
5.2 Experimental Setup for End-to-End Non-Verbal Speech Understaning
@@
In this task, we use Qwen2-Audio (Chu et al. 2024) as our base model. Qwen2-Audio is an advanced audio-language model capable of processing various audio inputs and generating either analytical results or textual responses. However, it can only classify non-verbal events or transcribe speech, and does not support integrating non-verbal tags into the transcription. To enable this capability, we introduce special tokens corresponding to non-verbal tags. We then fine-tune Qwen2-Audio on the CapSpeech (Wang et al. 2025) and NonVerbalSpeech-38K datasets to equip the model with the ability to understand and generate these tags. The details of the fine-tuning process can be found in the Appendix.
@@
To evaluate the model’s ability to integrate non-verbal tags into transcriptions, we conduct experiments on NonVerbalTTS (Borisov, Spirin, and Diatlova 2025), a high-quality, human-annotated real-world dataset specifically designed for non-verbal speech modeling. We select a subset of 754 samples containing only the three tags: [laughing], [breath], and [coughing]. For simplicity in evaluation, each utterance contains one tag. The number of samples per tag is listed in the Appendix.
In addition to the standard word error rate (WER), we use three tag-level metrics evaluating type and position accuracy: (1) F1-Score for tag correctness; and (2) Tag Position Distance (TPD) and its normalized version Normalized Tag Distance (NTD). TPD and NTD are computed after aligning predicted and reference tags using Levenshtein edit distance, where
$\text{TPD}=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-g_{i}|,\quad\text{NTD}=\frac{\text{TPD}}{L},$
with $p_{i}$, $g_{i}$ the positions of matched tags and $L$ the reference length. Lower TPD and NTD indicate better positional alignment. Detailed definitions are provided in the Appendix.
@@ 6.1 Experimental Results on Non-Verbal TTS
Table 4 summarizes non-verbal speech generation results for English (EN) and Chinese (ZH). Among baselines, CosyVoice2 excels in standard speech metrics (lowest WER/CER, highest DNSMOS and QMOS) but shows limited non-verbal controllability (lower CLAP-Score and IMOS). Dia performs best on English non-verbal metrics (CLAP and IMOS) but suffers from poor stability and high failure rates, especially unsupported in Chinese, leading to exclusion of its Chinese subjective scores. F5-TTS serves as a strong baseline with competitive performance metrics, achieving the highest SSIM. However, the pretrained F5-TTS does not support non-verbal TTS, and its CLAP-Score is reported for reference only. Our fine-tuned F5-TTS variant shows superior Chinese performance, achieving top subjective scores (QMOS, IMOS) and strong objective metrics, demonstrating competitive intelligibility and naturalness, as well as improved non-verbal controllability. CapSpeech and NVTTS underperform in Chinese due to English-only training data, reflected in lower QMOS, IMOS, SSIM and higher CER. On English tests, our NVS matches NVTTS on IMOS and CLAP-Score despite limited English data (4.6%). CapSpeech attains high CLAP but low DNSMOS and QMOS, likely due to synthetic non-verbal integration. When restricting the CLAP-Score evaluation to tags supported by CapSpeech (Table 5), as shown in the last line of Table 4, the model trained on the proposed dataset outperforms CapSpeech.
Overall, the diversity and naturalness of our dataset contribute to the competitive intelligibility, naturalness, and controllability achieved in multilingual non-verbal speech synthesis.
6.2 Experimental Results on End-to-End Non-Verbal Speech Understaning
Table 6 compares end-to-end non-verbal speech understanding performance between Whisper-Large-V3, Qwen2-Audio, and its fine-tuned variants. Whisper-Large-V3 achieves the best transcription accuracy (lowest WER). The original Qwen2-Audio shows much higher WER and does not support non-verbal tag integration, so tag-related metrics are unavailable. Fine-tuning Qwen2-Audio on our dataset significantly improves both transcription and non-verbal tag recognition. Our fine-tuned model attains higher F1-Score than the CapSpeech variant, indicating better tag detection accuracy. Although its TPD and NTD are slightly worse due to ASR transcription misalignment during tag insertion (Figure 2, Step 3), the results show that training on diverse real data improves robustness compared to CapSpeech, which relies on fixed sound event sets and limited speech corpora.
@@ 7 Discussion
Finally, we discuss the scalability of our pipeline, which is reflected in two main aspects:
  • The frame-level detection model is extensible to general sounds beyond non-verbal expressions by increasing the number of query tokens $N_{q}$. With CLAP (Elizalde et al. 2023) text embeddings as input, it also supports description-based sound and effect detection for finer-grained recognition.
  • Unlike prior works such as NonVerbalTTS (Borisov, Spirin, and Diatlova 2025), our pipeline is fully automated, requiring no manual verification. Compared to Capspeech (Wang et al. 2025), it yields more natural data distributions not constrained by reading-style corpora like LibriTTS (Zen et al. 2019).

### v2

@@ 2.2 Non-Verbal Segments Detection
CosyVoice2 $0.087$ $0.071$ $3.941$ $5.012$ 3.236{3.236} 3.280{3.280} $0.514$ $0.671$ $3.58_{(\pm 0.06)}$ $3.54_{(\pm 0.07)}$ $2.12_{(\pm 0.06)}$ $2.31_{(\pm 0.08)}$
Dia $0.182$ $0.097$ $15.687$ $171.882$ $2.749$ $2.528$ $0.431$ $0.319$ $3.47_{(\pm 0.13)}$ - $3.27_{(\pm 0.14)}$ -
F5-TTS - - $1.700$ $2.755$ $3.086$ $3.238$ $0.659$ $0.713$ $3.72_{(\pm 0.09)}$ $3.27_{(\pm 0.10)}$ - -
+ CapSpeech $0.125$ $0.121$ $2.791$ $9.856$ $3.009$ $3.153$ $0.570$ $0.669$ $3.66_{(\pm 0.10)}$ $2.40_{(\pm 0.07)}$ $1.79_{(\pm 0.07)}$ $1.88_{(\pm 0.08)}$
+ NVTTS $0.110$ $0.095$ $\mathbf{2.279}$ $7.042$ $3.099$ $\underline{3.248}$ $0.566$ $0.631$ $\mathbf{3.80}_{(\pm 0.08)}$ $3.28_{(\pm 0.08)}$ $22_{(\pm 0.10)} 3.3$$33_{(\pm 0.08)}$
+ NVSpeech $0.092$ $0.094$ $3.525$ $4.726$ $3.040$ $3.214$ $\mathbf{0.596}$ $\underline{0.683}$ $49_{(\pm 0.10)} 3.3$$32_{(\pm 0.07)}$ $2.83_{(\pm 0.11)}$ $3.36_{(\pm 0.09)}$
+ SMIIP-NV $0.257$ $0.281$ 3.501{3.501} 3.910{3.910} $\mathbf{3.132}$ $\mathbf{3.259}$ $0.497$ $0.604$ $75_{(\pm 0.08)} 3.3$$39_{(\pm 0.08)}$ $3.47_{(\pm 0.13)}$ $3.57_{(\pm 0.11)}$
+ SynParaSpeech $0.077$ $0.057$ $3.605$ $4.605$ $3.084$ $3.192$ $0.589$ $\mathbf{0.685}$ $3.25_{(\pm 0.08)}$ $3.00_{(\pm 0.07)}$ $2.88_{(\pm 0.12)}$ $3.12_{(\pm 0.09)}$
+ MNV-17 $0.087$ $0.100$ $5.457$ $4.670$ $3.102$ $\underline{3.248}$ $\underline{0.590}$ $0.667$ $3.43_{(\pm 0.11)}$ $\underline{3.41}_{(\pm 0.08)}$ $2.79_{(\pm 0.10)}$ $3.22_{(\pm 0.07)}$
[…]
+ NVS(TSA) $\mathbf{0.187}$ $\mathbf{0.179}$ $3.498$ $\mathbf{3.101}$ $3.114$ $3.236$ $0.567$ $0.670$ $\underline{3.76}_{(\pm 0.08)}$ $\mathbf{3.43}_{(\pm 0.09)}$ $\mathbf{3.48}_{(\pm 0.08)}$ $\mathbf{3.71}_{(\pm 0.07)}$
  • The best and second-best values in each column are highlighted in bold and underline, respectively. Note that CosyVoice2, Dia, and F5-TTS are excluded as reference models. Gray entries in CLAP Score and QMOS indicate that only a subset of NV types is supported and are therefore excluded from the ranking, see Table 3 for a more detailed comparison. NVTTS: NonVerbalTTS.
After training the frame-level detection model, we apply it to standardized waveforms using a sliding window strategy, as shown in Fig. 2 (Step 2). Standardized waveforms are used because tools like Silero-VAD often miss NV events at the beginning or end of speech segments. To provide sufficient context for detection, we use a window size of 10s and a hop size of 5s. To improve detection quality, we apply three filters: (1) discard segments shorter than 0.3 s; (2) remove segments with very low energy ($35\,\mathrm{dB}$), such as subtle breathing sounds; and (3) discard segments with a CLAP score below 0.3, since higher scores indicate true NV events. These thresholds are determined based on a simple inspection of the data.
2.3 Incorporating Non-Verbal Tags into Speech Content
After filtering out low-quality candidates, we match each NV segment with the nearest speech region obtained from preprocessing (Fig. 2(Step 3(1))). The pseudocode for this process is provided in the Appendix. It should be noted that NV events generally do not lie entirely within speech regions. Therefore, any NV segment located more than 1 second away from the nearest speech region is discarded.
After this, we obtain high-quality speech segments containing NV events, together with NV labels and their start/end timestamps, as well as speech transcriptions with word-level timestamps. To support NV speech generation and understanding, it is necessary to integrate the textual content with NV tags. As shown in Step 3 of Fig. 2, for a segment transcribed as “His funny face made us laugh,” a laughter sound occurs right after “His funny face.” By aligning their timestamps, we obtain the final transcript: “His funny face [laugh] made us laugh.” The algorithm is provided in the Appendix. However, the timestamp-based ordering (TBO) method assumes accurate word timestamps. In practice, ASR timestamps often contain offsets, causing misalignment between the TBO-generated NV-aware transcription and the speech. The Montreal Forced Aligner was expected to improve time stamp precision, but the resulting alignments show greater temporal drift.
To address this, we propose the Temporal-Semantic Alignment (TSA) method (Fig. 1, (c)), yielding two dataset variants (TBO and TSA). We first cut the speech at NV boundaries, run ASR on each segment, and merge the results with NV tags to obtain timing-accurate NV-aware transcriptions. Since this step may introduce semantic inconsistencies, we further align the NV-aware output with the full-segment ASR transcription using edit-distance alignment, combining accurate NV timings with coherent semantics. The final annotation integrates NV tags with the full-segment ASR output. NV events rarely overlap with semantic content, with <B> and </B> co-occurring in less than 2% of cases in the NVS dataset. Note that TSA does not rely on any word-level time stamps, which is a key advantage over TBO.
3 The NonVerbalSpeech-38K Dataset
Using the proposed NonVerbalSpeech-38K pipeline, we construct a dataset from diverse online speech sources (statistics are provided in the Appendix), covering a wide range of natural speaking styles and rich NVs. The initial release contains $\sim$131 hours of English and Chinese speech, totaling about 38,000 samples. As shown in Fig. 4, Chinese samples dominate due to the composition of the source data. Again, the frame-level detection model trained only on English still generalizes well to Chinese, indicating strong cross-lingual scalability. Frequent NV types such as [sigh], [sniff], [laughing], and [coughing] dominate, reflecting their prevalence in daily speech. Less frequent types (e.g., [crying], [yawn]) are also included. Further information and analysis can be found in the Appendix.
Figure 4: Language and NV label distribution in proposed NVS dataset.
Table 3: CLAP Score Differences (Comparison Systems - F5-TTS Fine-Tuned with NVS (TSA)), Broken Down by Non-Verbal Tag.
Model [laughing] [throat] [sigh] [coughing] [sniff] [breath]
English (EN)
CosyVoice2 0.132-0.132 - 0.149-0.149 0.150-0.150 - 0.056-0.056
Dia 0.050-0.050 0.034-0.034 $\mathbf{0.004}$ $\mathbf{0.062}$ 0.085-0.085 -
F5-TTS - - - - - -
+ CapSpeech 0.132-0.132 - - 0.151-0.151 - 0.015-0.015
+ NVTTS 0.160-0.160 0.003-0.003 0.136-0.136 0.043-0.043 0.043-0.043 0.077-0.077
+ NVSpeech 0.131-0.131 - 0.135-0.135 0.144-0.144 - 0.058-0.058
+ SMIIP-NV 0.072-0.072 - - $\mathbf{0.027}$ - -
+ SynParaSpeech 0.117-0.117 0.116-0.116 0.071-0.071 - - -
+ MNV-17 0.171-0.171 0.084-0.084 0.128-0.128 0.142-0.142 0.047-0.047 0.032-0.032
+ NVS(TBO) 0.076-0.076 0.043-0.043 0.105-0.105 0.099-0.099 0.074-0.074 0.051-0.051
Chinese (ZH)
CosyVoice2 0.134-0.134 - 0.124-0.124 0.152-0.152 - 0.098-0.098
Dia 0.105-0.105 $\mathbf{0.004}$ 0.056-0.056 0.132-0.132 0.150-0.150 -
F5-TTS - - - - - -
+ CapSpeech 0.157-0.157 - - 0.127-0.127 - 0.039-0.039
+ NVTTS 0.146-0.146 0.044-0.044 0.108-0.108 0.058-0.058 0.070-0.070 0.078-0.078
+ NVSpeech 0.140-0.140 - 0.103-0.103 0.118-0.118 - 0.055-0.055
+ SMIIP-NV 0.029-0.029 - - $\mathbf{0.058}$ - -
+ SynParaSpeech 0.120-0.120 0.113-0.113 0.060-0.060 - - -
+ MNV-17 0.180-0.180 0.052-0.052 0.092-0.092 0.105-0.105 0.015-0.015 0.033-0.033
+ NVS(TBO) 0.060-0.060 0.042-0.042 0.101-0.101 0.109-0.109 0.057-0.057 0.051-0.051
  • Negative values indicate that F5-TTS fine-tuned with NVS (TSA) outperforms the corresponding system. [throat] denotes [throatclearing], and “-” indicates unsupported NV types for the current comparison system.
4 Experiments
We evaluate the effectiveness of our dataset through two downstream tasks: NV speech generation and NV speech understanding. The details are as follows.
@@ 4.1 Experimental Setup for Non-Verbal Speech Generation
In the NV speech generation task, we use F5-TTS as the base model. To support NV control, we expand the vocabulary of F5-TTS by adding NV tags such as [breath], [sigh]. We fine-tune F5-TTS on seven datasets (Table 1), including the two variants of our NVS dataset. See Appendix for details.
@@
In addition to fine-tuning F5-TTS on various datasets, we include two strong baselines, CosyVoice2 (Du et al. 2024) and Dia (Nari-labs 2025), both supporting NV control and generation.
@@
Due to differences in NV support across models and datasets, we select six common NV phenomena for evaluation (Table 3). To evaluate NV speech generation, we use the prompt speech from the seed-tts-eval framework (Anastassiou et al. 2024). For text inputs, tagged sentences are generated using Gemini (Team et al. 2023) and GPT-4o (Hurst et al. 2024). For each NV tag, we construct 175 sentences in both Chinese and English, yielding 1,050 sentences per language. To evaluate system performance, we use standard metrics including DNSMOS, SSIM, and WER/CER to assess speech quality, speaker similarity, and intelligibility. Additionally, we introduce the CLAP Score to measure the presence of intended NV phenomena in synthesized speech. For subjective evaluation, 20 participants rated the samples on QMOS (for overall speech quality) and IMOS (for accuracy in rendering both verbal content and NV tags). See Appendix for details.
4.2 Experimental Setup for Non-Verbal Speech Understaning
@@
We use Qwen2-Audio and Whisper-Large-V3 as base models, neither of which supports NV-aware transcription. To enable this, we add special NV tokens and fine-tune the models on CapSpeech, NVSpeech, and both NVS variants. See Appendix for details.
@@
To evaluate the model’s ability to integrate NV tags into transcriptions, we use a subset of NonVerbalTTS as the English test set (754 samples, containing [laughing], [breath], and [coughing]), and a subset of SMIIP-NV and MNV-17 as the Chinese test set (784 samples, containing [laughing], [sigh], and [coughing]). All three datasets are manually constructed, ensuring high annotation quality (Table 1). In addition to standard WER/CER (excluding NV tags) and NV-aware WER/CER (NVWER/NVCER, including them), we use two tag-level metrics further evaluating position accuracy: Tag Position Distance (TPD) and its normalized version Normalized Tag Distance (NTD). TPD and NTD are computed after aligning predicted and reference transcriptions using Levenshtein edit distance, where
$\text{TPD}=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-g_{i}|,\quad\text{NTD}=\frac{\text{TPD}}{L},$ (4)
with $p_{i}$ and $g_{i}$ denoting matched tag positions in the prediction and reference and $L$ the reference length. Lower values indicate better accuracy. See Appendix for details.
@@ 5.1 Experimental Results on Non-Verbal TTS
Table 2 summarizes NV speech generation results for English and Chinese. Among baselines, CosyVoice2 excels in speech quality (highest DNSMOS and ZH QMOS) but shows limited NV controllability (low CLAP-Score and IMOS). Dia performs well on EN NV metrics (CLAP Score and IMOS) but suffers from poor stability and high failure rates, and is unsupported in Chinese, leading to the exclusion of its ZH subjective scores. F5-TTS is a strong base model, achieving the highest SSIM and lowest WER/CER. However, the pretrained F5-TTS does not support NV generation.
Fine-tuning on NVS (TSA) achieves superior Chinese performance, with top subjective scores, strong objective metrics, and the highest CLAP Score, indicating competitive intelligibility, naturalness, and NV controllability. CapSpeech and NVTTS underperform in ZH CER due to English-only training data. On English tests, NVS(TSA) also performs well, achieving the highest IMOS and CLAP Score, and second-highest DNSMOS and QMOS. CapSpeech attains a high CLAP Score but lowest DNSMOS and IMOS, likely due to unnatural synthetic non-verbal sounds. Tables 3 present CLAP Score comparisons per NV tag. We observe that NVS (TSA) ranks below existing models in only a few cases, demonstrating comprehensive and stable NV controllability. A similar trend is also observed for IMOS, see the Appendix for more details.
Overall, NVS (TSA) performs well across QMOS, IMOS, CLAP Score, ZH CER, and DNSMOS. Compared with NVS (TBO), it shows significant improvements in CLAP Score and IMOS, mainly because TSA provides more accurate annotations, particularly regarding NV positions, highlighting the importance of precise NV labels for controllable NV generation. Nonetheless, the NVS dataset exhibits some limitations in SSIM and English WER, likely due to its size, which we plan to address by further expanding the dataset in future work.
Table 4: EN Evaluation results for NV speech understanding.
Model WER (%) $\downarrow$ NVWER (%) $\downarrow$ TPD $\downarrow$ NTD $\downarrow$
Qwen2-Audio 33.087
+ CapSpeech 20.065 23.261 4.232 0.178
+ NVS(TBO) 16.901 21.952 6.223 0.291
+ NVS(TSA) 17.035 22.099 3.314 0.159
Whisper-Large-V3 17.542
+ CapSpeech 16.037 19.780 3.380 0.167
+ NVS(TBO) 15.469 21.843 8.490 0.396
+ NVS(TSA) 15.503 19.661 0.740 0.055
Table 5: ZH Evaluation results for NV speech understanding.
Model CER (%) $\downarrow$ NVCER (%) $\downarrow$ TPD $\downarrow$ NTD $\downarrow$
Qwen2-Audio 13.868
+ NVSpeech 4.314 4.994 1.530 0.050
+ NVS(TBO) 6.002 8.467 5.419 0.182
+ NVS(TSA) 4.921 6.093 1.737 0.062
Whisper-Large-V3 6.225
+ NVSpeech 4.841 5.166 0.904 0.030
+ NVS(TBO) 5.864 8.141 4.526 0.170
+ NVS(TSA) 5.087 5.844 1.110 0.038
5.2 Experimental Results on Non-Verbal Speech Understaning
Table 4 and 5 compares NV speech understanding performance. On the ZH test set, NVS (TSA) demonstrates competitive performance despite the smaller dataset size and fully automatic construction method without human effort (Table 1), with all metrics remaining low. On the EN test set, it consistently outperforms CapSpeech, which is constructed via rule-based synthesis and limited by dataset diversity. Notably, TSA shows significant improvements over TBO in TPD and NTD across both ZH and EN settings, further validating the superiority of the TSA method, consistent with Table 2.
@@ 6 Discussion
Finally, we highlight the scalability of our pipeline:
  • The unified NV detection model is language-agnostic, can generalize to other sounds by increasing query tokens $N_{q}$, and, with CLAP text embeddings, can also support description-based, fine-grained sound detection.
  • Unlike prior work that relies on human effort, our pipeline is fully automated, suggesting that the dataset can be further scaled—a direction for future work. Unlike rule-based synthesis methods, it captures natural NV sounds from diverse sources, which improves NV modeling.